2 meter of resistance wire of Cross section .5mm² has a resistance of 2.2ohm c calculate the resistivity of the meter

1
Expert's answer
2022-05-22T14:37:29-0400
"R=\rho\frac{l}{A}"

"\rho=\frac{RA}{l}\\
=\frac{2.2*0.5*10^{-6}}{2}=5.5*10^{-7}\:\rm \Omega\cdot m"